# Beyond Nash-Williams: Counterexamples to Clique Decomposition Thresholds for All Cliques Larger than Triangles

### v1

@@ 4.2. The case $q=4$
Let $v$ be an even integer such that $\alpha\cdot v$, $d_{2}\cdot v$, and therefore $d_{1}\cdot v$ are integers. Let $H$ be a $(d_{2}\cdot v)$-regular balanced bipartite graph on $v$ vertices, and let $I$ be an $(d_{1}\cdot v)$-regular balanced bipartite graph on $\alpha\cdot v$ vertices. Let $G$ be the join of $H$, $H$ and $I$.
@@
Recall that $H$ is bipartite and so, for every $F\in\binom{G}{K_{q}}$, we have $|V(F)\cap V(H)|\leq 2$. Thus we note that for every $\in\mathcal{F}_{0}, FF$ contains 22 $H$-internal edges and $4$ cross edges. For every $\in\mathcal{F}_{1}, FF$ contains 11 $H$-internal edge and $2$ cross edges (and $3$ small edges). For every $\in\mathcal{F}_{2}, FF$ contains 11 II-internal edges, at most $1$ cross edge, and at most 11 $H$-internal edge (and $4$ small edges). Therefore, similarly to the proof of Lemma˜4.2, we obtain, respectively from cross, $H$-internal and $I$-internal edges,
@@ 5. Acknowledgements
The first author would like to thank Victor Falgas Ravry and Maryam Sharifzadeh for assistance in locating the original thesis of Gustavsson.

### v2

@@ 4.2. The case $q=4$
Let $v$ be an even integer such that $\alpha\cdot v$, $d_{2}\cdot v$, and therefore $d_{1}\cdot v$ are even integers. Let $H$ be a $(d_{2}\cdot v)$-regular balanced bipartite graph on $v$ vertices, and let $I$ be an $(d_{1}\cdot v)$-regular balanced bipartite graph on $\alpha\cdot v$ vertices. Let $G$ be the join of $H$, $H$ and $I$.
@@
Recall that $H$ is bipartite and so, for every $F\in\binom{G}{K_{q}}$, we have $|V(F)\cap V(H)|\leq 2$. Thus we note that for every $\in\mathcal{F}_{0}, FF$ contains 22 $H$-internal edges and $4$ cross edges. For every $\in\mathcal{F}_{1}, FF$ contains 11 $H$-internal edge and $2$ cross edges (and $3$ small edges). For every $\in\mathcal{F}_{2}, FF$ contains 11 II-internal edges, at most $1$ cross edge, and at most 11 $H$-internal edge (and $4$ small edges). Therefore, similarly to the proof of Lemma˜4.1, we obtain, respectively from cross, $H$-internal and $I$-internal edges,
@@ 5. Acknowledgements
The authors would like to thank both referees for several helpful comments. The first author would like to thank Victor Falgas Ravry and Maryam Sharifzadeh for assistance in locating the original thesis of Gustavsson.